Circular No. 8492

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ao, 2005ap, 2005aq Three supernovae have been found on unfiltered CCD images: 2005ao

by T. Puckett and T. Crowley (cf. IAUC 8488; 0.50-m automated supernova patrol telescope); 2005ap by R. Quimby, P. Mondol, P. Hoeflich, J. C. Wheeler (all at the University of Texas), and C. Gerardy (Imperial College) using the 0.45-m ROTSE-IIIb telescope at McDonald Observatory; and 2005aq by J. Graham and W. Li (LOSS/KAIT; cf. IAUC 8484).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005ao Mar. 7.43 17^{h}44^{m}49.^{s}64 +61^{o}54′28′′.1 16.7 5′′.3 E, 10′′.9 S
2005ap Mar. 3.30 13 01 14.84 +27 43 31.4 18.5 3′′.5 W, 3′′.4 N
2005aq Mar. 7.13 4 31 38.82 − 4 35 06.8 17.3 1′′.7 E, 10′′.8 N

Approximate unfiltered CCD magnitudes, by the respective discoverers unless otherwise noted: SN 2005ao in NGC 6462, 2002 Feb. 13 UT, [20.0 (Puckett); 2003 Apr. 2, [20.0 (Puckett); 2005 Mar. 8.02, 16.6 (G. Sostero, Remanzacco, Italy). SN 2005ap, 2004 Dec. 15, [18.8; 2005 Jan. 14, [18.8; Mar. 4.29, 18.5. SN 2005aq in NGC 1599, Feb. 2.16, [18.0; Mar. 8.12, 17.2.

Quimby adds nothing is visible on the Digitized Sky Survey at the location of 2005ap, and a spectrum (range 430–890 nm), obtained on Mar. 7.46 with the 9.2-m Hobby/Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by B. Roman and V. Riley, shows a flat continuum with broad absorption at 526 and 557 nm; there are no features that would suggest a Galactic source. The spectrum is similar to the recent CfA spectrum of 2005an (cf. IAUC 8491), but of higher S/N; as with 2005an, Quimby et al. believe that SN 2005ap is an early type-II supernova, but given the general lack of distinct features, they cannot associate it with a template.

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2005ao, obtained on Mar. 8.52 UT by H. Hao with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova with a spectral-feature age (Riess et al. 1997, A.J. 114, 722) of ∼ −3 (± 2) days before maximum light. The supernova expansion velocity, derived from the minimum of Si II (rest 635.5 nm) and adopting the NED recession velocity of 11514 km/s for the host galaxy, is ∼ 11000 km/s. They add that a spectrogram (range 350–740 nm) of SN 2005aq, obtained on Mar. 8.11 UT, shows it to be a type-IIn supernova; the spectrum consists of strong, intermediate-width (FWHM 800 km/s) Balmer emission lines with broad wings that are superposed on a nearly featureless and flat continuum (assuming the NED recession velocity of 4042 km/s for the host galaxy).

2005 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green